A vector C has magnitude of 11 m/s and makes an angle 65° degree with the positive x-

axis, vector D has Cartesian co-ordinates Dx = 4m/s


2 and Dy = −3 m/s


2. Find the difference C⃗ -D�⃗ between


two vectors.


1
Expert's answer
2021-09-29T09:54:59-0400

Given:

"C_x=11*\cos65^{\circ}=4.65\:\rm m/s"

"C_y=11*\sin65^{\circ}=9.97\:\rm m/s"

"D_x=4.00\:\rm m/s"

"D_y=-3.00\:\rm m/s"



The difference between two vectors:


"{\bf C}-{\bf D}=(C_x-D_x){\hat i}+(C_y-D_y){\hat j}""=(4.65-4.00){\hat i}+(9.97-(-3.00)){\hat j}\\
=0.65{\hat i}+13.0{\hat j}"